OpenStudy (anonymous):

find domain of f(x) = 1+ e ^(-x)

OpenStudy (anonymous):

All real numbers. There's no restriction on what you can plug in for x

OpenStudy (anonymous):

can you show work for me ?

OpenStudy (anonymous):

There's not really any work, you just look at it and say i can put any x value in there and get a valid output...

OpenStudy (anonymous):

can you help me 1 more problem ?

OpenStudy (anonymous):

$3500 is invested at the rate of 9% compounded continuously . What is the balance at the end of 18 years ?

OpenStudy (anonymous):

I think you would use \[\large A=Pe^{rt}\] \[\huge A=3500e^{.09*18}\]

OpenStudy (anonymous):

thkss i got before you reply =))

OpenStudy (anonymous):

Yeah sorry had to look it up =/

OpenStudy (anonymous):

log (3x+7 ) + log ( x-2 )...solve x ???

OpenStudy (anonymous):

Theres no equal sign

OpenStudy (anonymous):

=1

OpenStudy (anonymous):

sorry my bad ... log (3x+7 ) + log ( x-2 ) =1

OpenStudy (anonymous):

ok well combine the left side \[\log [(3x+7)(x-2)]=1\] then apply base 10 exp. to both sides which will cancel the log \[(3x+7)(x-2)=10^1\]

OpenStudy (anonymous):

Now what?

OpenStudy (anonymous):

log = 10 ?

OpenStudy (anonymous):

No...theres no more log we got rid of it, solve for x

OpenStudy (anonymous):

ok ok thks so much , give you 10 medals

OpenStudy (anonymous):

determine which of the following angles of t = 2pi/7